# Revisiting the Orbital Dynamics of the Hot Jupiter WASP–12b with New Transit Times

### v1

@@ 4 Transit Light-curve Fitting
    Only complete and high-quality transit light curves (LCs) were included.
@@
Each of the WASP-12b transit light curves was then individually loaded into TAP to determine its mid-transit time.
To initiate the transit LC analysis, we fixed the orbital eccentricity ($e$) and longitude of periastron ($\omega$) to zero, following J. A. Carter & J. N. Winn (2009) and F. Fressin et al. (2010). For the remaining system parameters, we adopted the approach of S. Biswas et al. (2024, 2025), allowing the orbital period ($P$), scaled semi-major axis ($a/R_{\star}$), orbital inclination ($i$), and limb-darkening coefficients ($u_{1},u_{2}$) to vary within Gaussian priors constrained to $\pm 1\sigma$ of literature values. The planet-to-star radius ratio ($R_{p}/R_{\star}$) and the mid-transit time ($T_{m}$) were treated as completely free parameters.
For the TESS light curves of WASP-12b, the initial values of the linear and quadratic limb-darkening coefficients ($u_{1},u_{2}$) were estimated by interpolating the stellar parameters, effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), metallicity ([Fe/H]), and microturbulent velocity ($V_{t}$), from the tables of A. Claret (2017). For all other ground-based transit light curves observed in the clear, $V$, II, and $R$ filters, as well as the Sloan $g$, $r$, and $z$ filters, we followed L.-H. Su et al. (2021) and linearly interpolated the coefficients from the tables of A. Claret & S. Bloemen (2011) using the EXOFAST package (J. Eastman et al. 2013). The adopted stellar parameters were $T_{\mathrm{eff}}=6265$ K, $\log g=4.11$ cm s-2, and [Fe/H] = 0.12, as reported by P. Leonardi et al. (2024).
Since the clear filter covers both the $V$ and $R$ bands (G. Maciejewski et al. 2013), the limb-darkening coefficients for WASP-12b LCs obtained with this filter were calculated as the average of the $V$- and $R$-band values. For light curves observed with the Luminance filter, the coefficients derived for the $V$ filter were adopted. For the dataset of G. Maciejewski et al. (2013), we directly used the $u_{1}$ and $u_{2}$ values reported in their work. For the 23 light curves from K. A. Collins et al. (2017), we adopted their procedure, using coefficients calculated for the $R$ and Kepler passbands to represent their Clear with Blue Block (CBB; a high-pass filter with a cutoff near 500 nm) and open filters, respectively. For G. Maciejewski et al. (2016), we adopted the values provided for the Clear filter as representative for the None filter.
The initial parameter values were set as follows: $a/R_{\star}$, $i$, and $R_{p}/R_{\star}$ from I. Wong et al. (2022); orbital period ($P$) from P. Leonardi et al. (2024); and $T_{m}$, which was initialized automatically by TAP for WASP-12b. These initial parameter settings are listed in Table 6, while the interpolated limb-darkening coefficients for different filters are summarized in Table 7.
@@
Parameter Initial Value During MCMC Chain
P (days) 1.0914210a A Gaussian Prior with $\sigma$ = 0.00000020a
i (deg) 83.54b A Gaussian Prior with $\sigma$ = 0.74b
$a/R_{\ast}$ 3.061b A Gaussian Prior with $\sigma$ = 0.034b
$R_{p}/R_{\ast}$} 0.11600b Free
$T_{m}$} Set by eye Free
$u_{1}$ According to filter A Gaussian Prior with $\sigma$ = 0.05c
$u_{2}$ According to filter A Gaussian Prior with $\sigma$ = 0.05c
@@
Notes: a The initial and the prior value of the parameter P is directly adopted from P. Leonardi et al. (2024).
a The initial values of the parameters i, $a/R_{\ast}$, and $R_{p}/R_{\ast}$}, and the gaussian priors for i, $a/R_{\ast}$, are directly adopted from I. Wong et al. (2022).
c The priors of $u_{1}$ and $u_{2}$ taken from A. Claret (2017).
Table 7: The Calculated Values for Theoretical Limb-darkening Coefficients
Object Name Filter $u_{1}$ $u_{2}$
WASP-12 V 0.5224 0.2193
R 0.4181 0.2510
I 0.3288 0.2552
Clear 0.4702 0.2352
Luminance 0.5224 0.2193
TESS 0.5353 0.3456
@@
For each fitted parameter, the 50th percentile of the posterior probability distribution was adopted as the best-fit value, and the 15.9th and 84.1st percentiles were taken as the lower and upper $1\sigma$ uncertainties (i.e., the 68% credible intervals). All derived mid-transit times were consistent with previously reported values at the $1\sigma$ level. Graphical representations of the newly acquired normalized light curves from various ground-based observatories are shown in Figures 1, 2, and 3. while those from TESS, ETD, and ExoClock are presented in Figures 5-9, 10-11, and 14-15, respectively.
Figure 1: Light curve of WASP-12b observed by the 1.3 m Devasthal Fast Optical Telescope at Devasthal, Nainital of Aryabhatta Research Institute of Observational Sciences (ARIES). Here, the error bars corresponding to the normalized flux are shown in light blue.
Figure 2: Light curves of WASP-12b observed by the 0.61 m telescope installed at IERCOO campus of ICSP, Kolkata. Here, the error bars corresponding to the normalized flux are shown in purple.
Figure 3: Light curves of WASP-12b observed by the 0.3 m AG Optical IDK telescope at the Utah Desert Remote Observatory in Beryl, Utah. Here, the error bars corresponding to the normalized flux are shown in light green.
5 Timing Analysis
Before starting the transit timing analysis, to avoid redundancy, we excluded 7 duplicate mid-transit times reported in both the ETD and ExoClock databases, retaining the mid-transit times from the ETD as the preferred values, as the ETD data are more consistent and of higher quality (Data Quality (DQ) index below 3). Additionally, the ETD light curves have less noise and smaller uncertainties than the ones from ExoClock. To support our decision, we included a comparison table 8 that shows the transit times reported by both ETD and ExoClock at same epoch, along with their uncertainties. This comparison confirms that the ETD data are consistent with the ExoClock data, further supporting the reliability of the ETD dataset.
@@ 5 Timing Analysis
After applying the above mentioned correction, we modeled the transit timing data using three distinct approaches, following the methodologies described in K. C. Patra et al. (2017) and S. W. Yee et al. (2020). In this analysis, we combined all 391 mid-transit times obtained from both space-based and ground-based observations.The first model is the standard constant-period assumption, represented by a linear ephemeris. The linear ephemeris is expressed as,
Table 8: Comparison of 7 mid-transit times obtained from ETD and ExoClock at identical epochs.
Observer Name Date of Observation Filter ETD Timings Exoclock Timings
Anaël Wünsche 17-01-19 clear $2458501.3926\pm 0.0027$ $2458501.3924\pm 0.0030$
Anaël Wünsche 21-02-21 R $2458536.3167\pm 0.0017$ $2458536.3168\pm 0.0019$
Anaël Wünsche 22-02-19 R $2458537.4063\pm 0.0017$ $2458537.4081\pm 0.0021$
Ferran Grau Horta 11-01-20 R $2458860.4699\pm 0.0011$ $2458860.4695\pm 0.0012$
Thomas Mollier 10-12-20 clear $2459194.4429\pm 0.0016$ $2459194.4406\pm 0.0027$
Ferran Grau Horta 19-02-21 R $2459265.38422\pm 0.00078$ $2459265.38395\pm 0.00080$
Fabio Salvaggio 15-12-21 R $2459564.4278\pm 0.0021$ $2459564.4335\pm 0.0023$
@@
where $E$ is the transit epoch, $T_{0}$ is the reference mid-transit time at epoch $E=0$ (taken as the first transit of WASP-12b observed by L. Hebb et al. 2009), $P$ is the orbital period, and $T_{c}(E)$ is the calculated mid-transit time at epoch $E$.
Then we calculated the timing residuals, O-C (pronounced “O minus C”; C. Sterken 2005), where O denotes the observed mid-transit times and C is the calculated mid-transit times. In the absence of transit timing variation, we would anticipate no significant deviations of the derived O-C (observed-minus-calculated) values from zero. However, we observed a significant deviation on both sides of zero and this deviation from the zero value could be the first indication of timing anomalies. The estimated timing residuals (O-C) along with their corresponding epochs and original mid-transit times ($T_{m}$) are shown in Table LABEL:tab:6.
@@
0 2454515.52548 -0.0030324 L. Hebb et al. (2009) This Paper
13 2454529.71660 -0.0003577 L. Hebb et al. (2009) This Paper
@@
References. L. Hebb et al. (2009); T. Chan et al. (2011); G. Maciejewski et al. (2011, 2013, 2016); K. A. Collins et al. (2017); S. W. Yee et al. (2020); E. R. Adams et al. (2024); E. Alvarado et al. (2024).
@@
By substituting the derived values of $P$ and $\frac{dP}{dE}$ (see Table 10) for WASP-12b into equation (4) of K. C. Patra et al. (2017),
@@
we obtain the orbital decay rate as $\dot{P}\approx-30.31\pm 0.92~\mathrm{ms~yr^{-1}}$. This result provides strong evidence that the orbit of WASP-12b is decaying rapidly, as inferred from the full set of available transit-timing measurements spanning a baseline of $\sim 15$ years.
The third model considered is the apsidal precession model, which assumes that the planet’s orbit is slightly eccentric and that the argument of periastron ($\omega$) undergoes uniform precession. The analytical form of this model, originally proposed by A. Giménez & M. Bastero (1995), is given by:
@@
where the free parameters are defined as follows: $T_{ap0}$ is the mid-transit time at $E=0$, $P_{s}$ is the sidereal period, $e$ is the orbital eccentricity, $\omega_{0}$ is the argument of periastron at the reference epoch ($E=0$), and $\frac{d\omega}{dE}$ is the precession rate of the periastron. In this context, $E$ denotes the epoch, $T_{ap}(E)$ represents the calculated mid-transit time at epoch $E$, and $\omega$ corresponds to the angle between the ascending node in the plane of the sky and the orbital periastron.
@@
By analyzing the timing residuals obtained after fitting the three different timing models, we construct the observed minus calculated (O-C) diagram, as shown in Figure  4. The O-C diagram serves as a powerful diagnostic tool for detecting long-term variations in the orbital period. In this case, the orbital decay model, represented by the red dashed line, reveals a clear downward trend. Furthermore, we randomly drew a sample of 100 parameter sets from the posterior distributions of the orbital decay model (represented by the brown solid lines in the O-C diagram) and extrapolated them over the next $\sim$13 years to investigate the projected evolution of the decay. The resulting curves consistently exhibit the same declining trend, confirming the robustness of the predicted orbital decay behavior. In the O-C diagram, the blue dashed curve represents the apsidal precession model. To investigate the future behavior of WASP-12b under apsidal precession, we applied the same procedure as for the orbital decay model. We randomly drew a sample of 100 parameter sets from the posterior distributions of the apsidal precession model (represented by the cyan line) and extrapolated them over the next $\sim$13 years. The resulting O-C diagram indicates that the apsidal precession model does not exhibit any significant deviation from the linear ephemeris.
@@
Figure 4: O-C diagram for analysing 391 mid-transit times of WASP-12b. The blue filled square show the data from E. R. Adams et al. (2024), the red filled squares are from T. Chan et al. (2011), the green filled square is from K. A. Collins et al. (2017), the black filled circle is from E. Alvarado et al. (2024), the brown filled asterisks are from L. Hebb et al. (2009), the magenta filled circles are from P. Leonardi et al. (2024), the yellow filled upward triangles are from G. Maciejewski et al. (2011), the green filled circles are from G. Maciejewski et al. (2013), the orange filled right triangles are from the quality 1 and quality 2 data of ETD, the blue filled circles are from Exoclock, the red filled circle is from G. Maciejewski et al. (2016), the blue filled asterisk is from S. W. Yee et al. (2020), the yellow filled squares are from TESS data, the purple filled squares are from the new ground-based data for 7 nights. The dashed black line, red and blue curves represent the linear, orbital decay and apsidal precession models. The lines are drawn for 100 randomly chosen sets of parameters from the Markov chains of posteriors of the orbital decay (brown) and apsidal precession (cyan) models. The models are extrapolated for the next $\sim 10$ years to illustrate the broad spectrum of possible solutions.
@@
Mid-transit time T0 BJDTDB 2454515.52851242 $^{+0.528\times 10^{-4}}_{-0.528\times 10^{-4}}$
@@
Mid-transit time Tq0 BJDTDB 2454515.52535025 $^{+0.11\times 10^{-3}}_{-0.11\times 10^{-3}}$
Decay Rate dP/dE days/epoch -1.05 $\times$ 10-9 $^{+0.32\times 10^{-10}}_{-0.32\times 10^{-10}}$
Decay Rate dP/dt ms/yr -30.31 0.92
@@
Sidereal Period Ps days 1.091418841 $^{+2.43\times 10^{-8}}_{-2.10\times 10^{-8}}$
Mid-transit time Tap0 BJDTDB 2454515.52859484 $^{+5.77\times 10^{-5}}_{-6.39\times 10^{-5}}$
Eccentricity e 0.0031 $^{+3.27\times 10^{-5}}_{-4.76\times 10^{-5}}$
Argument of Periastron $\omega$0 rad 3.11 $^{+0.026}_{-0.105}$
Precession Rate $\omega/d$E rad/epoch 0.0010 $^{+1.53\times 10^{-5}}_{-3.73\times 10^{-5}}$
@@
$\chi^{2}$, $\chi^{2}_{red}$ 1520.84, 3.94
AIC 1531.08
BIC 1550.92
5.1 Goodness-of-fit metrics
To assess the goodness of fit, we calculated the reduced chi-square value ($\chi^{2}$ per degree of freedom, denoted as $\chi^{2}_{r}$) for the three best-fit models (Table 10) using the expression ${\chi}^{2}_{r}$ = ${\chi}^{2}$/n, where $n$ is the number of degrees of freedom. Among the tested models, the orbital decay model yields the lowest reduced chi-square value ($\chi^{2}_{r}=1.55$ with 388 degrees of freedom) compared to the linear model ($\chi^{2}_{r}=4.31$ with 389 degrees of freedom) and the apsidal precession model ($\chi^{2}_{r}=3.94$ with 386 degrees of freedom). This indicates that the orbital-decay model provides a superior fit to the transit timing data of WASP-12b.
To further evaluate the statistical preference among the models, we employed two widely used information criteria: the Akaike Information Criterion (AIC; H. Akaike 1974) and the Bayesian Information Criterion (BIC; G. Schwarz 1978). These are defined as $\mathrm{AIC}=\chi^{2}+2k_{F}$, $\mathrm{BIC}=\chi^{2}+k_{F}\ln N_{P}$, where $k_{F}$ is the number of free parameters in the model, and $N_{P}$ is the total number of data points ($N_{P}=391$ in this analysis). For the models considered here, $k_{F}=2$ for the linear ephemeris, $k_{F}=3$ for the orbital decay model, and $k_{F}=5$ for the apsidal precession model. The corresponding values of $\chi^{2}$, AIC, and BIC are reported in Table 10.

### v2

@@ IV Transit Light-curve Fitting
    Only complete and high-quality transit light curves were included.
@@
Each WASP-12 b transit light curve was analyzed individually using juliet in order to determine the corresponding mid-transit time. In our fits, we sampled the transformed parameters r1_p1r1\_p1 and r2_p1r2\_p1, which parameterize the planet-to-star radius ratio, $p=R_{\rm p}/R_{\ast}$, and the impact parameter, defined as $b=(a/R_{\ast})\cos i$. We also fitted the limb-darkening parametrization $q_{1,\mathrm{TESS}}$ and $q_{2,\mathrm{TESS}}$, which parameterize the quadratic limb-darkening coefficients $u_{1}$ and $u_{2}$ following the quadratic law of Kipping (2013). The transformation from the $(r_{1},r_{2})$ plane to the $(b,p)$ plane was implemented using the formalism described by Espinoza (2018).
For the TESS light curves, the initial values of quadratic limb-darkening coefficients ($u_{1}$, $u_{2}$) were obtained by interpolating the stellar effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), metallicity ($\mathrm{Fe/H}]$), and microturbulent velocity ($V_{\mathrm{t}}$) from the limb-darkening tables of Claret (2017). For all other ground-based transit light curves observed in the clear, $V$, II, luminance, and $R$ filters, as well as the Sloan $r$ and $z$ bands, we followed the procedure of Su et al. (2021) and interpolated the quaratic limb-darkening coefficients from the tables of Claret and Bloemen (2011) using the EXOFAST package (Eastman et al. 2013).
The adopted stellar parameters were $T_{\mathrm{eff}}=6265\,\mathrm{K}$, $\log g=4.11\,\mathrm{cm\,s^{-2}}$, and $\mathrm{Fe/H}]=0.12$, as reported by Leonardi et al. (2024b). Since the clear filter spans both the $V$ and $R$ passbands (Maciejewski et al. 2013b, the limb-darkening coefficients for light curves obtained in this filter were taken as the average of the corresponding $V$ and $R$ band values. For light curves observed with the luminance filter, the coefficients derived for the $V$ band were adopted.
For the dataset of Maciejewski et al. (2013b), the limb-darkening coefficients ($u_{1}$, $u_{2}$) reported in their study were used directly. For the 23 light curves from Collins et al. (2017), we followed their methodology and adopted coefficients calculated for the $R$ and Kepler passbands to represent their Clear with Blue Block (CBB; a high-pass filter with a cutoff near 500 nm) and open filters, respectively. For Maciejewski et al. (2016b), the coefficients provided for the Clear filter were adopted as representative for the None filter.
The prior distributions adopted for the planetary and instrumental parameters are listed in Table 6, while the parameterized limb-darkening coefficients ($q_{1}$ and $q_{2}$) for the different filters are summarized in Table 7.
@@
Name of Parameters Description Prior
Planetary Parameters
P (days) Orbital period $\mathcal{N}(1.0914210,\;0.00000020)^{a,\dagger}$
[…]
r1_p1 Transformed transit parameter (jointly parameterizes $b$ and $R_{\mathrm{p}}/R_{\star}$) $\mathcal{U}(0.37215959921,0.39005188)$
r2_p1 Transformed transit parameter (jointly parameterizes $b$ and $R_{\mathrm{p}}/R_{\star}$) $\mathcal{U}(0,1)$
e Eccentricity of the orbit $\mathrm{Fixed}\,(0.0)$
$\omega$ (degree) Argument of periastron $\mathrm{Fixed}\,(90\degree.0)$
Instrumental parameters
$\rho$ Mean stellar density $\mathcal{N}(0,\,0.1)$
mdilution Dilution factor $\mathrm{Fixed}\,(1.0)$
mflux Relative flux offset $\mathcal{N}(0,\,0.1)$
sigma_w jitter $\mathcal{N}(0,\,0.1)$
$\sigma_{GP}$ Amplitude of the GP $\log\mathcal{U}(10^{-6},\,10^{6})^{\dagger}$
$\rho_{\mathrm{GP}}$ Time/length-scale of the Matern part of the GP $\log\mathcal{U}(10^{-3},\,10^{3})$
$q_{1}$ Parameterized limb-darkening coefficient $\mathcal{N}($According to filter, $0.05)^{b}$
$q_{2}$ Parameterized limb-darkening coefficient $\mathcal{N}($According to filter, $0.05)^{b}$
@@
Notes: a The initial and the prior value of the parameter P is directly adopted from Leonardi et al. (2024a).
b The priors of $q_{1}$ and $q_{2}$ taken from Claret (2017).
$\mathcal{N}(\mu,\sigma)$ represents a normal prior, $\mathcal{U}(a,b)$ a uniform prior, and $\log\mathcal{U}(a,b)$ a log-uniform prior.
§ Here, tmid1 and tmid2 are the lower and upper bounds for mid-transit time for each light curve.
* r1_p1 and r2_p1 are the reparameterized transit parameters following Espinoza et al. (2019). The quadratic limb-darkening coefficients are parameterized as $(q_{1},q_{2})$ following Kipping (2013).
Table 7: The Calculated Values for parameterized Limb-darkening Coefficients
Filter $q_{1}$ $q_{2}$
V 0.4690 0.2783
R 0.3741 0.2395
I 0.2808 0.2072
Clear 0.4202 0.2600
Sloan r 0.4016 0.2481
Sloan z 0.2393 0.1906
Luminance 0.4690 0.2783
TESS 0.4560 0.3293
@@
For each fitted parameter, the 50th percentile of the posterior probability distribution was adopted as the best-fit value, while the 15.9th and 84.1st percentiles were taken as the lower and upper $1\sigma$ uncertainties, corresponding to the 68% credible intervals. Graphical representation of the newly acquired 7 normalized light curves from various ground-based observatories is shown in Figure 1. While those from TESS, ETD, and ExoClock are presented in Figures 3-7, 8-9, and 12-13, respectively.
Figure 1: Seven new light curves of WASP-12 b observed with the 1.3 m Devasthal Fast Optical Telescope (top), 0.61 m VASISTHA (middle), and the 0.3 m IDK telescope (bottom). The upper panels show the normalized light curves, while the lower panels display the corresponding residuals. The data shown have been corrected by subtracting the best-fitting Gaussian Process (GP) model and the out-of-transit (OOT) linear baseline function. The solid line represents the best-fitting transit model obtained through juliet. Flux uncertainties are indicated by error bars, shown in red, green, and purple for the three telescopes, respectively.
V Timing Analysis
Before initiating the transit-timing analysis, we identified seven transits that are reported in both the ETD and ExoClock databases. They are obviously the same light curves as the observer names, telescopes, and observatory information are identical, in addition, the observation dates and filters are also the same. Observers submitted the same light curves to both ETD and ExoClock databases. We verified that the corresponding mid-transit times from the two databases are consistent within their quoted uncertainties. To avoid double-counting the same observational epochs, we retained a single representative mid-transit time per transit in the final timing analysis.
In these cases, we adopted the mid-transit times derived from our fits to the ETD light curves as the preferred values, since the corresponding light curves exhibit superior quality, characterized by lower noise levels, smaller timing uncertainties, and Data Quality (DQ) indices below 3.
@@ V Timing Analysis
After applying the above mentioned correction, we modeled the transit timing data using three distinct approaches, following the methodologies described in Patra et al. (2017) and Yee et al. (2020). In this analysis, we combined all 391 mid-transit times obtained from both space-based and ground-based observations.The first model is the standard constant-period assumption, represented by a linear ephemeris. The linear ephemeris is expressed as,
@@
where $E$ is the transit epoch, $T_{0}$ is the reference mid-transit time at epoch $E=0$ (taken as the first transit of WASP-12 b observed by Hebb et al. 2009), $P$ is the orbital period, and $T_{c}(E)$ is the calculated mid-transit time at epoch $E$.
To refine the best-fit parameters obtained from linear fitting, we adopted a Gaussian likelihood with uniform priors on $P$ and $T_{0}$ and sampled the posterior using the affine-invariant MCMC algorithm of Goodman and Weare (2010). Convergence and sampling efficiency were assessed using the mean acceptance fraction ($a_{f}$), integrated autocorrelation time ($\tau$), and effective sample size ($N_{\rm eff}$). The chains exhibit robust convergence, with $a_{f}\simeq 0.44$ (within the optimal range 0.2–0.5) and $N_{\rm eff}>50$ per walker, satisfying standard MCMC criteria (Foreman-Mackey et al., 2013; Stefansson et al., 2017). MCMC non-convergence can lead to unstable model behavior (Hagey et al., 2022). We therefore evaluated convergence using the integrated autocorrelation time ($\tau$) (Goodman and Weare, 2010; Foreman-Mackey et al., 2013), obtaining $\tau\simeq 19$ steps. The effective number of independent samples is $N_{\rm eff}\simeq 1053$, exceeding the recommended minimum threshold of 50 per walker, as advised by the emcee framework. Then we calculated the timing residuals, O-C (pronounced “O minus C”; Sterken 2005), where O denotes the observed mid-transit times and C is the calculated mid-transit times. In the absence of transit timing variation, we would anticipate no significant deviations of the derived O-C (observed-minus-calculated) values from zero. However, we observed a significant deviation on both sides of zero and this deviation from the zero value could be the first indication of timing anomalies. The estimated timing residuals (O-C) along with their corresponding epochs and original mid-transit times ($T_{m}$) are shown in Table LABEL:tab:6.
@@
0 2454515.52548 -0.0033471 Hebb et al. (2009) This Paper
13 2454529.71660 -0.0006704 Hebb et al. (2009) This Paper
@@
References. Hebb et al. (2009); Chan et al. (2011); Maciejewski et al. (2011b, 2013a, 2016a); Collins et al. (2017); Yee et al. (2020); Adams et al. (2024); Alvarado et al. (2024).
@@
For the orbital-decay model, we followed the same fitting procedure as for the linear model, but ran 32 000 MCMC steps per walker to sample the posterior distribution and determine the best-fit ephemeris and uncertainties for the three free parameters: $P$, $T_{0}$, and $\mathrm{d}P/\mathrm{d}E$. The initial 60 steps (approximately twice the integrated autocorrelation time) were discarded as burn-in to mitigate the effects of correlated samples (Almenara et al., 2016). The prior on the period derivative, $\mathrm{d}P/\mathrm{d}E$, was allowed to vary freely over both positive and negative values.
By substituting the derived values of $P$ and $\frac{dP}{dE}$ (see Table 9) for WASP-12 b into equation (4) of Patra et al. (2017),
@@
we obtain the orbital decay rate as $\dot{P}\approx-31.97\pm 0.80~\mathrm{ms~yr^{-1}}$. This result provides strong evidence that the orbit of WASP-12 b is decaying rapidly, as inferred from the full set of available transit-timing measurements spanning a baseline of $\sim 15$ years.
The third model considered is the apsidal precession model, which assumes that the planet’s orbit is slightly eccentric and that the argument of periastron ($\omega$) undergoes uniform precession. The analytical form of this model, originally proposed by Giménez and Bastero (1995), is given by:
@@
where the free parameters are defined as follows: $T_{ap0}$ is the mid-transit time at $E=0$, $P_{s}$ is the sidereal period, $e$ is the orbital eccentricity, $\omega_{0}$ is the argument of periastron at the reference epoch ($E=0$), and $\frac{d\omega}{dE}$ is the precession rate of the periastron. In this context, $E$ denotes the epoch, $T_{ap}(E)$ represents the calculated mid-transit time at epoch $E$, and $\omega$ corresponds to the angle between the ascending node in the plane of the sky and the orbital periastron. To infer the best-fit ephemeris of the apsidal precession model, we followed the same fitting approach outlined for linear model fitting and orbital decay model fitting, while extending the MCMC sampling to $10^{5}$ steps per walker.
@@
The best-fitting ephemerides derived from fitting the linear, orbital-decay, and apsidal-precession models are summarized in Table 9. By analyzing the timing residuals obtained after fitting the three different timing models, we construct the observed minus calculated (O-C) diagram, as shown in Figure  2. The O-C diagram serves as a powerful diagnostic tool for detecting long-term variations in the orbital period. In this case, the orbital decay model, represented by the red dashed line, reveals a clear downward trend. Furthermore, we randomly drew a sample of 100 parameter sets from the posterior distributions of the orbital decay model (represented by the brown solid lines in the O-C diagram) and extrapolated them over the next $\sim$13 years to investigate the projected evolution of the decay. The resulting curves consistently exhibit the same declining trend, confirming the robustness of the predicted orbital decay behavior. In the O-C diagram, the blue dashed curve represents the apsidal precession model. To investigate the future behavior of WASP-12 b under apsidal precession, we applied the same procedure as for the orbital decay model. We randomly drew a sample of 100 parameter sets from the posterior distributions of the apsidal precession model (represented by the cyan line) and extrapolated them over the next $\sim$13 years. The resulting O-C diagram indicates that the apsidal precession model does not exhibit any significant deviation from the linear ephemeris.
@@
Figure 2: O-C diagram for analysing 391 mid-transit times of WASP-12b. The blue filled square show the data from Adams et al. (2024), the red filled squares are from Chan et al. (2011), the green filled square is from Collins et al. (2017), the black filled circle is from Alvarado et al. (2024), the brown filled asterisks are from Hebb et al. (2009), the magenta filled circles are from Leonardi et al. (2024b), the yellow filled upward triangles are from Maciejewski et al. (2011a), the green filled circles are from Maciejewski et al. (2013a), the orange filled right triangles are from the quality 1 and quality 2 data of ETD, the blue filled circles are from Exoclock, the red filled circle is from Maciejewski et al. (2016a), the blue filled asterisk is from Yee et al. (2020), the yellow filled squares are from TESS data, the purple filled squares are from the new ground-based data for 7 nights. The dashed black line, red and blue curves represent the linear, orbital decay and apsidal precession models. The lines are drawn for 100 randomly chosen sets of parameters from the Markov chains of posteriors of the orbital decay (brown) and apsidal precession (cyan) models. The models are extrapolated for the next $\sim 10$ years to illustrate the broad spectrum of possible solutions.
@@
Mid-transit time T0 BJDTDB 2454515.52882715 $^{+0.494\times 10^{-4}}_{-0.495\times 10^{-4}}$
@@
Mid-transit time Tq0 BJDTDB 2454515.52526934 $^{+0.10\times 10^{-3}}_{-0.10\times 10^{-3}}$
Decay Rate dP/dE days/epoch -1.10577512 $\times$ 10-9 $^{+0.28\times 10^{-10}}_{-0.28\times 10^{-10}}$
Decay Rate dP/dt ms/yr -31.97 0.80
@@
Sidereal Period Ps days 1.09141868 $^{+3.64\times 10^{-8}}_{-2.17\times 10^{-8}}$
Mid-transit time Tap0 BJDTDB 2454515.52891730 $^{+6.19\times 10^{-5}}_{-7.48\times 10^{-5}}$
Eccentricity e 0.0030 $^{+3.24\times 10^{-5}}_{-1.75\times 10^{-3}}$
Argument of Periastron $\omega$0 rad 3.10 $^{+0.037}_{-3.56}$
Precession Rate $\omega/d$E rad/epoch 0.0011 $^{+0.59\times 10^{-5}}_{-1.78\times 10^{-5}}$
@@
$\chi^{2}$, $\chi^{2}_{red}$ 2183.76, 5.66
AIC 2194.10
BIC 2213.94
V.1 Goodness-of-fit metrics
To assess the goodness of fit, we calculated the reduced chi-square value ($\chi^{2}$ per degree of freedom, denoted as $\chi^{2}_{r}$) for the three best-fit models (Table 9) using the expression ${\chi}^{2}_{r}$ = ${\chi}^{2}$/n, where $n$ is the number of degrees of freedom. Among the tested models, the orbital decay model yields the lowest reduced chi-square value ($\chi^{2}_{r}=2.23$ with 388 degrees of freedom) compared to the linear model ($\chi^{2}_{r}=6.25$ with 389 degrees of freedom) and the apsidal precession model ($\chi^{2}_{r}=5.66$ with 386 degrees of freedom). This indicates that the orbital-decay model provides a superior fit to the transit timing data of WASP-12 b.
To further evaluate the statistical preference among the models, we employed two widely used information criteria: the Akaike Information Criterion (AIC; Akaike 1974) and the Bayesian Information Criterion (BIC; Schwarz 1978). These are defined as $\mathrm{AIC}=\chi^{2}+2k_{F}$, $\mathrm{BIC}=\chi^{2}+k_{F}\ln N_{P}$, where $k_{F}$ is the number of free parameters in the model, and $N_{P}$ is the total number of data points ($N_{P}=391$ in this analysis). For the models considered here, $k_{F}=2$ for the linear ephemeris, $k_{F}=3$ for the orbital decay model, and $k_{F}=5$ for the apsidal precession model. The corresponding values of $\chi^{2}$, AIC, and BIC are reported in Table 9.